OpenStudy (anonymous):

a current of 4 A flows from a battery when a light bulb is connected across its terminals. The P.D across the terminal is 12V. a) what quantity of charge leaves the battery every second. b) how much potential energy does each coulomb leaving the battery possess. c) how much charge must be given out by the battery if it is to supply 60 joules of energy? d) How long does the battery take to supply 60 joules of energy?

OpenStudy (souvik):

current is 4 A that means 4 C charge per second

OpenStudy (anonymous):

please also send the reply of b) c) and d)

OpenStudy (souvik):

the each coulomb possess the energy E=qV=(1C)*(12V)=12 J

OpenStudy (souvik):

energy E=QV...E=60J and V=12V ...you will get Q...

OpenStudy (souvik):

current I=Q/t....I=4 A and you know Q from the above equation...you will get time t...

OpenStudy (souvik):

if any problem...ask me

OpenStudy (anonymous):

thanks

OpenStudy (souvik):

welcome..:)

OpenStudy (anonymous):

in 10s, a charge of 25c leaves a battery and 200 j of energy are delivered to an outside circuit as result. a) what is P.D across the battery? b) what current flows from the battery?

OpenStudy (souvik):

current=charge/time...so what is the current here?

OpenStudy (souvik):

and energy E=Q*V...you will get V...

OpenStudy (souvik):

tell me the answers?

OpenStudy (anonymous):

8 volt and current is 2.5 amp

OpenStudy (souvik):

good job!!!:)

OpenStudy (anonymous):

thanks

OpenStudy (souvik):

close it then...

OpenStudy (anonymous):

wht?

OpenStudy (souvik):

the question

OpenStudy (anonymous):

wht is difference between voltage and current

OpenStudy (souvik):

it may help you http://hyperphysics.phy-astr.gsu.edu/hbase/emcon.html#emcon